# DietGlance: Dietary Monitoring and Personalized Analysis at a Glance with Knowledge-Empowered AI Assistant

### v2

@@ 4.3.3 Nutritional Analysis and Dietary Suggestion
With the RAG-grounded LLM, we first use a crafted prompt (available in Supplement 1.3) to derive nutritional analysis based on the description and amount of food and drinks, including the nutrient and corresponding values. The total nutrition consumed per meal was calculated from the estimated portions. The system then checked if each nutrient intake was too high, too low, or within a reasonable range, offering suggested reference values.
@@
We further fed the meal logs, nutrition analysis, and user profile into the RAG-grounded LLM to generate general dietary suggestions and personalized suggestions tailored to users’ dietary goals, with references for these suggestions. User profiles could include general demographic information (i.e., gender, age, height, weight) and personalized dietary goals, such as muscle development, weight loss, weight control, allergen, and vegetarianism. The meal logs contained the description of meals including the consumed food and drink and the meal timing. The system used this comprehensive information to generate dietary suggestions using the RAG-grounded LLM. As shown in Fig. 6 (yellow part), a query containing the information of the user profile, meal logs, nutritional analysis, and the dietary goal is first used to retrieve the relevant contexts from the nutrition library. The contexts and the query will be used to build the prompt for producing tailored dietary suggestions leveraging the contextual and reasoning capabilities of the LLM. For those without specific goals, the system provided balanced diet suggestions based on their profile and dietary behavior.
@@
Fig. 7: Interface Overview. There are desktop and mobile versions. The interface includes four main parts: (A) Meal Image Logging, (B) Interactive Nutritional Analysis, (C) General and Personalized Dietary Suggestions, and (D) Context-Aware AI Chatbot.
4.4 Interface Design
@@
To support users exploring the results, seeking further details, such as explanations for specific nutrients, alternative food options, or personalized meal plans, and providing contextual information, we developed an interactive desktop and mobile interface (Fig. 7) that not only displays results and suggestions but also allows users to correct errors, add missing details, and access more information. The interface consists of four major components: (1) Meal Image Logging (Fig. 7-A), (2) Interactive Nutritional Analysis (Fig.7-B), (3) Dietary Suggestion (Fig. 7-C), and (4) Context-Aware AI Chatbot (Fig. 7-D). Chronologically displayed meal log thumbnails allowed users to review their dietary patterns and access privacy-protected meal images (Fig. 7-A, R1, R4), while clicking on a meal session provided detailed nutritional analysis and total nutrient consumption (Fig.7-B, R1, R2). The interactive analysis enabled users to edit meal summaries and dynamically update nutritional insights and suggestions. Personalized dietary suggestions (Fig.7-C, R3), limited to seven items for clarity, offered general guidance for a balanced diet and tailored suggestions based on user goals, with transparent source references to enhance trust. The AI chatbot (Fig.7-D, R3) provided context-aware guidance, leveraging user profiles, meal logs, and prior interactions to deliver accurate, personalized responses, further streamlining the experience with preloaded common questions. This combination of features supported reflective, informed, and personalized dietary management. More details are available in the Supplement 2.3.
@@
5 Study I: Short-Term Study for Evaluating DietGlance from the four design requirements
@@
5.1 Study I.1: Short-Term User Study to Explore Users’ Perceptions towards DietGlance (R1-R4)
@@
We implemented the proposed system and conducted a user study recruiting participants to collect data using Aria Glasses in their daily routine and explore the dietary monitoring analysis results via the user interface. Through this study, we aimed to evaluate the system’s performance on the four design requirements qualitatively and quantitatively.
5.1.1 Participants and Data Collection
@@
We recruited 33 participants (P1-P33), including 21 self-declared males (63.6%) and 12 self-declared females (36.4%), aged 23 to 73 ($M=30.73$, $SD=10.51$). The group consisted of twenty-one university students, nine senior researchers, one accountant, one retired person, and one engineer. Twenty-one (63.6%) participants self-reported exp. with dietary monitoring and analysis, while the remaining participants (36.4%) were inexp. More details are available in Supplement 3.
@@
Participants used Aria Glasses before, during, and after their meal sessions, capturing the ground truth video recordings, audio, and IMU data. Due to the power limitation of Aria Glasses, we did not ask the participants to wear the Aria Glasses all day but for periods that covered the meal sessions, which also included various behaviors like walking, sitting, talking, using mobile phones, etc. Participants reviewed their recordings and labeled the start and end of each ingestive episode, verified by the authors. Participants earned an average wage of about \$8 per hour of recording and would get about \$6.42 after finishing the user study additionally.
5.1.2 Data Processing and Ingestive Episode Detection Performance
@@
Recordings were segmented into one-second windows, and if ingestive behavior occurred for more than 50% of the window, it was labeled as ingestive; otherwise, it was labeled as others (non-eating or drinking). In total, we collected 38.23 hours of recordings, covering 144 meal sessions. Each participant contributed between two (one day without breakfast) and nine sessions ($M=4.4$, $SD$ = 1.74). Of the total, 24.78 hours were ingestive episodes, and 13.46 hours were other behaviors. The performance of ingestive episode detection was evaluated using the leave-one-user-out (LOUO) approach, our system achieved an F1-score of 0.925, with precision at 0.939 and recall at 0.912. The LOUO method evaluates the model’s generalization capacity by iteratively training it on data from all users except one, testing on the excluded user’s data, and repeating this process for each user, thereby offering a comprehensive, user-specific assessment of model performance. Despite some time windows being misclassified, the system successfully captured the meal images across the entire meal sessions. Additionally, during image segmentation, images without target items were excluded, further enhancing system robustness and ensuring successful image capture for all 144 meal sessions.
5.1.3 Study Settings
@@
After obtaining participants’ consent, we introduced the system and provided a brief tutorial on the user interface. Participants were then asked to explore their dietary monitoring results, review meal identifications for accuracy, examine dietary suggestions and nutritional analysis, and use the chatbot. We recorded their screens and tracked interactions using Google Analytics [35]. Participants then completed a survey, including the System Usability Scale (SUS) [36], Net Promoter Score (NPS) [69], User Experience Questionnaire (UEQ) [68], and questions corresponding to the four design requirements. Finally, semi-structured interviews were conducted for open-ended feedback on their experience. Survey scales were converted to a 5-point Likert scale (1 = strongly disagree, 5 = strongly agree) to make it consistent. Detailed scale items and interview scripts are in Supplement 4. The lead author first performed a deductive thematic analysis [12] using predefined themes derived from the system’s design requirements for the interviews and textual data from participants, and iteratively discussed with other authors to validate the codes and address any disagreements.
5.1.4 User Study Results
@@
Our evaluation revealed broadly positive usability, satisfaction, and user experience across both exp. and inexp. participants. Exp. participants consistently rated the system higher in terms of ease of use, automation, and efficiency, while inexp. participants highlighted challenges in adapting to workflows and interacting with the system. Satisfaction scores averaged 8.3 (NPS), with high utility acknowledged by most participants. However, the physical discomfort from prolonged wearing of glasses (such as heat buildup around the glasses’ legs, sweat accumulation, and slippage due to the limited glasses sizes provided) and workflow complexity emerged as areas for improvement, particularly for less tech-savvy users. Details on these assessments, including SUS, NPS, and UEQ metrics, are provided in Supplement 5.
@@
Participants spent an average of 10.10 minutes per session (2.59 minutes per meal). Reviewing dietary suggestions averaged 2.82 minutes. Inspecting nutritional analysis averaged 5.05 minutes, indicating participants prioritized understanding their diet over simply viewing suggestions. Manual corrections were minimal, with 27 changes across 144 meal sessions, mostly involving beverages (e.g., from “Coke” to “Coke Zero”). Users also corrected misidentified items or cooking methods (e.g., “grilled pork” to “grilled duck” or “steamed” to “fried”). Despite features designed to reduce errors of including nearby meals, occasional deletions of extra items indicate room for improvement. Participants also engaged actively with the chatbot, logging 125 conversation pairs (average of 3.79 per participant), primarily for meal planning with constraints such as “in Chinese style.” Iterative interactions, like P6 refining meal plans to exclude breakfast or avoid certain foods (full chat history is available in Supplement 6), demonstrated the chatbot’s adaptability in addressing preferences and fostering personalized dietary management. Overall, the system’s core features were actively utilized, with strong engagement enhancing the personalized user experience and demonstrating the system’s value in dietary monitoring and analysis.
@@
5.2 Study I.2: Quantitative Evaluation with Crowd Workers on Diet Identification and Logging Performance (R1)
@@
5.2.1 Evaluation Settings
We used a crowdsourcing approach, presenting processed meal images (extracted from the video but cropped for privacy) alongside the LLM-generated text descriptions to crowd workers recruited from AWS MTurk [15]. Crowd workers worked independently. Specifically, due to the complexity of dishes and the high burdensome of measuring exact ingredient quantities, we employed an indirect evaluation of diet item quantity estimation by evaluating the nutrient estimation performance, since nutrient estimation involves quantity inference, as detailed in Sec. 5.3.
5.2.2 Evaluation Metrics
The crowd workers were tasked with determining the following:
  • Correct Items ($C$): Items in the ground truth images correctly identified in the text description.
@@ 5.2.2 Evaluation Metrics
Each meal was labeled by 10 qualified crowd workers. Workers could get \$0.1 for each qualified label. We then calculated precision, recall, and F1-score.
5.2.3 Results
Based on the crowd worker labels, the proposed system achieved an average F1-score of 0.972, with a precision of 0.957 and a recall of 0.989, demonstrating the effectiveness of DietGlance in diet identification. The meal that achieved the lowest precision (0.643) included a bowl of noodles with various ingredients, such as blood cubes, bean sprouts, sliced meat, green onions, and tomatoes, making it difficult for LLM to correctly identify the ingredients from the processed meal log images. The meal with the lowest recall (0.853) contained a bowl of rice with grilled fish, lemonade (with small pieces of onion and Seaweed on the top), a bowl of clear soup, and a cup of milk tea. DietGlance failed to describe those small ingredients. Overall, the system performed well, particularly on Western-style dishes with clearly separated components. Highly complex meals like Malatang and Hot Pot did not perform the worst; this might be because their ingredients are very difficult to describe even for the crowd workers.
5.3 Study I.3: Quantitative Evaluation with Experts on Nutritional Analysis Performance (R2)
We held a panel with Food and Nutrition Science experts who highlighted the difficulty of determining nutrient values due to factors like ingredient freshness, cooking methods, portion sizes, storage conditions, and ingredient interactions during cooking [82], etc. Despite this variability, experts can estimate nutrients using food composition databases, standardized recipes, and their experience with similar meals.
5.3.1 Evaluation Settings
@@
5.3.2 Evaluation Metrics
We measured expert agreement using the Intraclass Correlation Coefficient (ICC), showing high reliability of the average estimation from experts with ICC(2,k) > 0.9 (more details in Supplement 7). Therefore, to evaluate DietGlance’s performance, we compared its estimation to the expert averages using Mean Absolute Percentage Error (MAPE) due to varying nutrient units and scales, calculated as: ${MAPE=\frac{1}{N}\sum_{n=1}^{N}|\frac{\overline{E}_{n}-S_{n}}{\overline{E}_{n}}|\times 100\%}$, where $N$ is the number of meals. $\overline{E}_{n}$ and $S_{n}$ are the average estimations of experts and the system, respectively.
Fig. 8: The results of MAPE (%) in Studies I.3 and I.4. Lower values indicate better performance.
5.3.3 Results
As shown in Fig. 8, the system demonstrated high accuracy for key macronutrients, including energy (9.13%), protein (11.70%), total fat (12.80%), and carbohydrates (10.05%). These macronutrients are major food components and are typically well-documented in databases, making them easier for DietGlance to estimate accurately. Given the challenges in nutrient estimation, such as variations in food composition databases, portion size estimation errors, and the difficulty of accurately assessing mixed dishes, a MAPE around 10% is considered a strong result [32, 77].
Moderate accuracy was achieved for nutrients like calcium (25.19%), iron (24.76%), cholesterol (22.75%), dietary fiber (21.27%), and zinc (25.28%), possibly due to their greater variability across food sources and bioavailability, complicating precise estimation. The system exhibited poor alignment with experts’ estimation of phosphorus (53.02%) and potassium (74.90%). These nutrients are known to exhibit significant variability across different foods due to a variety of factors, such as plant and animal growth conditions, differences in food type, preparation, and source [57]. This variability, combined with the effects of processing and cooking methods, makes accurate estimation particularly challenging.

### v3

@@ 4.3.3. Nutritional Analysis and Dietary Suggestion
With the RAG-grounded LLM, we first use a crafted prompt (available in Appendix B.3) to derive nutritional analysis based on the description and amount of food and drinks, including the nutrient and corresponding values. The total nutrition consumed per meal was calculated from the estimated portions. The system then checked if each nutrient intake was too high, too low, or within a reasonable range, offering suggested reference values.
@@
We further fed the meal logs, nutrition analysis, and user profile into the RAG-grounded LLM to generate general dietary suggestions and personalized suggestions tailored to users’ dietary goals, with references for these suggestions. User profiles could include general demographic information (i.e., gender, age, height, weight) and personalized dietary goals, such as muscle development, weight loss, weight control, allergens, and vegetarianism. The meal logs contained the description of meals, including the consumed food and drink, and the meal timing. The system used this comprehensive information to generate dietary suggestions using the RAG-grounded LLM. As shown in Fig. 6 (yellow part), a query containing the information of the user profile, meal logs, nutritional analysis, and the dietary goal is first used to retrieve the relevant contexts from the nutrition library. The contexts and the query will be used to build the prompt for producing tailored dietary suggestions leveraging the contextual and reasoning capabilities of the LLM. For those without specific goals, the system provided balanced diet suggestions based on their profile and dietary behavior.
@@
Fig. 7. Interface Overview. There are desktop and mobile versions. The interface includes four main parts: (A) Meal Image Logging, (B) Interactive Nutritional Analysis, (C) General and Personalized Dietary Suggestions, and (D) Context-Aware AI Chatbot.
4.4. Interface Design
@@
To support users exploring the results, seeking further details, such as explanations for specific nutrients, alternative food options, or personalized meal plans, and providing contextual information, we developed an interactive desktop and mobile interface (Fig. 7) that not only displays results and suggestions but also allows users to correct errors, add missing details, and access more information. The interface consists of four major components: (1) Meal Image Logging (Fig. 7-A), (2) Interactive Nutritional Analysis (Fig.7-B), (3) Dietary Suggestion (Fig. 7-C), and (4) Context-Aware AI Chatbot (Fig. 7-D). Chronologically displayed meal log thumbnails allowed users to review their dietary patterns and access privacy-protected meal images (Fig. 7-A, R1, R4), while clicking on a meal session provided detailed nutritional analysis and total nutrient consumption (Fig.7-B, R1, R2). The interactive analysis enabled users to edit meal summaries and dynamically update nutritional insights and suggestions. Personalized dietary suggestions (Fig.7-C, R3), limited to seven items for clarity, offered general guidance for a balanced diet and tailored suggestions based on user goals, with transparent source references to enhance trust. The AI chatbot (Fig.7-D, R3) provided context-aware guidance, leveraging user profiles, meal logs, and prior interactions to deliver accurate, personalized responses, further streamlining the experience with preloaded common questions. This combination of features supported reflective, informed, and personalized dietary management. More details are available in the Appendix C.3.
@@
5. Study I: Short-Term Study for Evaluating DietGlance from the four design requirements
@@
5.1. Study I.1: Short-Term User Study to Explore Users’ Perceptions towards DietGlance (R1-R4)
@@
We implemented the proposed system and conducted a user study, recruiting participants to collect data using Aria Glasses in their daily routine and explore the dietary monitoring analysis results via the user interface. Through this study, we aimed to evaluate the system’s performance on the four design requirements qualitatively and quantitatively.
5.1.1. Participants and Data Collection
@@
We recruited 33 participants (P1-P33), including 21 self-declared males (63.6%) and 12 self-declared females (36.4%), aged 23 to 73 ($M=30.73$, $SD=10.51$). The group consisted of twenty-one university students, nine senior researchers, one accountant, one retired person, and one engineer. Twenty-one (63.6%) participants self-reported exp. with dietary monitoring and analysis, while the remaining participants (36.4%) were inexp. More details are available in Appendix D.
@@
Participants used Aria Glasses before, during, and after their meal sessions, capturing the ground truth video recordings, audio, and IMU data. Due to the power limitation of Aria Glasses, we did not ask the participants to wear the Aria Glasses all day, but for periods that covered the meal sessions, which also included various behaviors like walking, sitting, talking, using mobile phones, etc. Participants reviewed their recordings and labeled the start and end of each ingestive episode, verified by the authors. Participants earned an average wage of about \$8 per hour of recording and would get about \$6.42 after finishing the user study.
5.1.2. Data Processing and Ingestive Episode Detection Performance
@@
Recordings were segmented into one-second windows, and if ingestive behavior occurred for more than 50% of the window, it was labeled as ingestive; otherwise, it was labeled as others (non-eating or drinking). In total, we collected 38.23 hours of recordings, covering 144 meal sessions ($M=15.9$ minutes, Min=3.3=3.3, Max=66.2=66.2, $SD=15.9$). Each participant contributed between two (one day without breakfast) and nine sessions ($M=4.4$, SDSD = 1.74, spanning one to three days). Of the total, 24.78 hours were ingestive episodes, and 13.46 hours were other behaviors. The performance of ingestive episode detection was evaluated using the leave-one-user-out (LOUO) approach. Our system achieved an F1-score of 0.925, with precision at 0.939 and recall at 0.912. The LOUO method evaluates the model’s generalization capacity by iteratively training it on data from all users except one, testing on the excluded user’s data, and repeating this process for each user, thereby offering a comprehensive, user-specific assessment of model performance. Despite some time windows being misclassified, the system successfully captured the meal images across the entire meal sessions. Additionally, during image segmentation, images without target items were excluded, further enhancing system robustness and ensuring successful image capture for all 144 meal sessions.
5.1.3. Study Settings
@@
After obtaining participants’ consent, we introduced the system and provided a brief tutorial on the user interface. Participants were then asked to explore their dietary monitoring results, review meal identifications for accuracy, examine dietary suggestions and nutritional analysis, and use the chatbot. We recorded their screens and tracked interactions using Google Analytics (Ledford et al., 2011). Participants then completed a survey, including the System Usability Scale (SUS) (Lewis, 2018), Net Promoter Score (NPS) (Score, 2018), User Experience Questionnaire (UEQ) (Schrepp et al., 2014), and questions corresponding to the four design requirements. Finally, semi-structured interviews were conducted for open-ended feedback on their experience. Survey scales were converted to a 5-point Likert scale (1 = strongly disagree, 5 = strongly agree) to make it consistent. Detailed scale items and interview scripts are in Appendix E. The lead author first performed a deductive thematic analysis (Bowman et al., 2023) using predefined themes derived from the system’s design requirements for the interviews and textual data from participants, and iteratively discussed with other authors to validate the codes and address any disagreements.
5.1.4. User Study Results
@@
Our evaluation revealed broadly positive usability, satisfaction, and user experience across both exp. and inexp. participants. Exp. participants consistently rated the system higher in terms of ease of use, automation, and efficiency, while inexp. participants highlighted challenges in adapting to workflows and interacting with the system. Satisfaction scores averaged 8.3 (NPS), with high utility acknowledged by most participants. However, the physical discomfort from prolonged wearing of glasses (such as heat buildup around the glasses’ legs, sweat accumulation, and slippage due to the limited glasses sizes provided) and workflow complexity emerged as areas for improvement, particularly for less tech-savvy users. Details on these assessments, including SUS, NPS, and UEQ metrics, are provided in Appendix F.
@@
Participants spent an average of 10.10 minutes per session (2.59 minutes per meal). Reviewing dietary suggestions averaged 2.82 minutes. Inspecting nutritional analysis averaged 5.05 minutes, indicating participants prioritized understanding their diet over simply viewing suggestions. Manual corrections were minimal, with 27 changes across 144 meal sessions, mostly involving beverages (e.g., from “Coke” to “Coke Zero”). Users also corrected misidentified items or cooking methods (e.g., “grilled pork” to “grilled duck” or “steamed” to “fried”). Despite features designed to reduce errors of including nearby meals, occasional deletions of extra items indicate room for improvement. Participants also engaged actively with the chatbot, logging 125 conversation pairs (average of 3.79 per participant), primarily for meal planning with constraints such as “in Chinese style.” Iterative interactions, like P6 refining meal plans to exclude breakfast or avoid certain foods (full chat history is available in Appendix G), demonstrated the chatbot’s adaptability in addressing preferences and fostering personalized dietary management. Overall, the system’s core features were actively utilized, with strong engagement enhancing the personalized user experience and demonstrating the system’s value in dietary monitoring and analysis.
@@
5.2. Study I.2: Quantitative Evaluation with Crowd Workers on Diet Identification and Logging Performance (R1)
@@
5.2.1. Evaluation Settings
We used a crowdsourcing approach, presenting processed meal images (extracted from the video but cropped for privacy) alongside the LLM-generated text descriptions to crowd workers recruited from AWS MTurk (Cheung et al., 2017) who met specific qualification criteria i.e., Human Intelligence Task (HIT) Approval Rate $\geq 90\%$ and Approved HITs $\geq 500$. Crowd workers worked independently. Specifically, due to the complexity of dishes and the high burden of measuring exact ingredient quantities, we employed an indirect evaluation of diet item quantity estimation by evaluating the nutrient estimation performance, since nutrient estimation involves quantity inference, as detailed in Sec. 5.3.
5.2.2. Evaluation Metrics
The crowd workers were tasked with verifying the alignment between the images and the generated text by counting the following:
  • Correct Identified Items ($C$): Items in the ground truth images correctly identified in the text description.
@@ 5.2.2. Evaluation Metrics
Submissions failing these quality checks were automatically re-released to new workers to maintain the target sample size. Each meal was labeled by 10 qualified crowd workers. Workers could get \$0.1 for each qualified label. We then calculated precision, recall, and F1-score.
5.2.3. Results
Based on the crowd worker labels, the proposed system achieved an average F1-score of 0.972, with a precision of 0.957 and a recall of 0.989, demonstrating the effectiveness of DietGlance in diet identification. The meal that achieved the lowest precision (0.643) included a bowl of noodles with various ingredients, such as blood cubes, bean sprouts, sliced meat, green onions, and tomatoes, making it difficult for LLM to correctly identify the ingredients from the processed meal log images. The meal with the lowest recall (0.853) contained a bowl of rice with grilled fish, lemonade (with small pieces of onion and Seaweed on the top), a bowl of clear soup, and a cup of milk tea. DietGlance failed to describe those small ingredients. Overall, the system performed well, particularly on Western-style dishes with clearly separated components. Highly complex meals like Malatang and Hot Pot did not yield the lowest F1-scores; this might be because their ingredients are very difficult to describe even for the crowd workers.
5.3. Study I.3: Quantitative Evaluation with Experts on Nutritional Analysis Performance (R2)
We held a panel with Food and Nutrition Science experts who highlighted the difficulty of determining nutrient values due to factors like ingredient freshness, cooking methods, portion sizes, storage conditions, and ingredient interactions during cooking (Willett, 2012), etc. Despite this variability, experts can estimate nutrients using food composition databases, standardized recipes, and their experience with similar meals.
5.3.1. Evaluation Settings
@@
5.3.2. Evaluation Metrics
We measured expert agreement using the Intraclass Correlation Coefficient (ICC), showing high reliability of the average estimation from experts with ICC(2,k) ¿ 0.9 (more details in Appendix H). Therefore, to evaluate DietGlance’s performance, we compared its estimation to the expert averages using Mean Absolute Percentage Error (MAPE) due to varying nutrient units and scales, calculated as: $MAPE=\frac{1}{N}\sum_{n=1}^{N}|\frac{\overline{E}_{n}-S_{n}}{\overline{E}_{n}}|\times 100\%$}, where $N$ is the number of meals. $\overline{E}_{n}$ and $S_{n}$ are the average estimations of experts and the system, respectively.
Fig. 8. The results of MAPE (%) in Studies I.3 and I.4. Lower values indicate better performance.
5.3.3. Results
As shown in Fig. 8, the system demonstrated high accuracy for key macronutrients, including energy (9.13%), protein (11.70%), total fat (12.80%), and carbohydrates (10.05%). These macronutrients are major food components and are typically well-documented in databases, making them easier for DietGlance to estimate accurately. Given the challenges in nutrient estimation, such as variations in food composition databases, portion size estimation errors, and the difficulty of accurately assessing mixed dishes, a MAPE around 10% is considered a strong result (Kipnis et al., 2002; Thames et al., 2021).
Moderate accuracy was achieved for nutrients like calcium (25.19%), iron (24.76%), cholesterol (22.75%), dietary fiber (21.27%), and zinc (25.28%), possibly due to their greater variability across food sources and bioavailability, complicating precise estimation. The system exhibited poor alignment with experts’ estimation of phosphorus (53.02%) and potassium (74.90%). These nutrients are known to exhibit significant variability across different foods due to a variety of factors, such as plant and animal growth conditions, differences in food type, preparation, and source (Organization and others, 2004). This variability, combined with the effects of processing and cooking methods, makes accurate estimation particularly challenging.